40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 3 18:42:43 2014, Jenne, Update, LSC, Green PDH box boosts BoostCartoon.pdf
    Reply  Thu Sep 4 00:56:44 2014, Jenne, Update, LSC, Green PDH box boosts TFstage_newDesign_3Sept2014.pngBoostNoBoost_newDesign_3Sept2014.pngBoostOnly_newDesign_3Sept2014.pngD0901351-v1_3Sept2014.pdf
       Reply  Thu Sep 4 03:12:55 2014, ericq, Update, LSC, Green PDH box boosts boostedTF.pdfgain.jpg
       Reply  Thu Sep 4 16:45:10 2014, Jenne, Update, LSC, Green PDH box boosts IMG_1712.JPGIMG_1714.JPG
          Reply  Fri Sep 5 05:32:57 2014, Jenne, Update, LSC, Green PDH box boosts TRX_TRY_ALSvsIR_4Sept2014.pdf
             Reply  Fri Sep 5 21:13:57 2014, Jenne, Update, LSC, Green PDH out of loop ALS_XY_outOfLoop_5Sept2014.pdf
Message ID: 10450     Entry time: Thu Sep 4 03:12:55 2014     In reply to: 10448
Author: ericq 
Type: Update 
Category: LSC 
Subject: Green PDH box boosts 

Jenne made her board modifications, and the measured TF agreed with the design. Alas, the green would not lock to the arm in this state. 

I think that the reason is that the new TF does not have nearly as much low frequency gain as the old one, for a given UGF. Thus, for example, the 1Hz noise due to the pendulum resonance, has 30dB less loop gain suppressing it. 

boostedTF.pdf

 

NEED MORE gain.jpg

 

ELOG V3.1.3-